# Automating Android Build Repair: Bridging the Reasoning-Execution Gap in LLM Agents with Domain-Specific Tools

### v2

@@ 1 Introduction
Android is the largest and most active mobile development platform in the world, with a market share of 73.9% of the global mobile OS market, over 2.2 million apps on the Google Play Store, and over 1.4 million questions tagged "android" on Stack Overflow (statista_googleplay_2024; statcounter_os_share; stackoverflow_android_questions). Despite its popularity, a practical challenge remains: applications cannot reliably be built. A comprehensive study of 5,222 apps revealed that only 31.3% could be built successfully out-of-the-box (liu2024understanding). In our dataset, we find the primary source of issues is syntax errors (Table 3(a)). While Continuous Integration and Continuous Deployment (CI/CD) pipelines are a well-established practice for mitigating these errors (rostami2023usage; baitha2024streamlining; ghaleb2024ci; liu2022first), liu2024understanding reported only 21.8% out of the 5,222 apps had CI/CD scripts. Furthermore, even with CI/CD, maintaining a stable build requires significant effort. hassan2017automatic found that over 74% of CI/CD configuration changes are dedicated to build fixing and environment updates. This maintenance is necessary to address a complex range of errors, including syntax mistakes, missing resource files, configuration errors, and library unavailability.
Large Language Models (LLMs) and their application in Automated Program Repair (APR) offer a promising alternative (zhang2024systematic; yin2024thinkrepair; xia2023automated; jin2023inferfix). However, the application of LLMs to Android build errors remains an underexplored area. Existing LLM coding benchmarks, like SWE-bench (jimenez2023swe), primarily evaluate models on resolving bugs from GitHub issues, not the transient, environment-specific build failures common in Android development. Furthermore, existing datasets of Android build failures (liu2024understanding; hassan2017automatic) capture a repository at a single, broken point in time, making fixes difficult to verify as feasible without a corresponding working build to serve as ground truth.
@@
Second, we propose GradleFixer, an agent-based approach for build error fixing using domain-specific tools. Inspired by prior work showing that specialized tools improve agent performance (singh2025code; wang2023voyager) (details in Appendix A), GradleFixer equips an LLM agent with domain-specific tools, which are simple wrappers for shell commands, to inspect and manipulate the Gradle build environment. Our experiments show that GradleFixer significantly outperforms state-of-the-art baselines across all categories of build failures (Figure 1). We infer from GradleFixer’s success that LLMs have the high-level knowledge needed to fix build errors. The poorer performance of shell-based agents, however, suggests this knowledge fails to transfer to effective low-level execution. This gap explains why other effective agents also rely on specialized toolsets (wang2023voyager; singh2025code), although these studies do not explore this choice as a research question. Our approach, which we term Tool Bridging, addresses this gap by replacing a general-purpose shell with domain-aware abstractions to bridge high-level reasoning to correct low-level execution. We hypothesize this strategy improves performance through two simultaneous mechanisms. First, by providing API-like tools that abstract away complex command syntax, it allows the model to focus on reasoning rather than implementation. Second, it constrains the action space, reducing irrelevant operations. By reframing the task into a format the model uses more reliably (liu2024toolace; zhang2024ecoact; zhang2025nemotron), our strategy connects the LLM’s high-level reasoning to effective low-level execution. Our contributions are summarized as follows:
@@ 2.1 Project Curation
AndroidBuildBench is curated from 43 active Android projects on GitHub. These were selected by filtering the top 100 most-starred applications for Java/Kotlin usage, >500 stars, >100 closed pull requests, recent commits, and a Gradle build script. This process yielded 43 open-source projects (Appendix G). A summary of the project statistics is shown in Table 1. We focus on the Gradle build system, as it is used by 95.1% of Android applications according to a recent study (liu2024understanding).
@@ 2.3 Error Categorization
We manually analyzed the 184 test set issues, adopting the five build failure categories proposed by liu2024understanding.
@@
Our analyses are presented in Table 3(a). A notable finding is the absence of NDK errors. We posit this absence results from our curation method, which is anchored to successfully built pull requests. NDK errors often result from build machine environment incompatibilities. Since our method requires a successful build as an endpoint, it naturally filters out states with systemic environment issues. The errors captured are those transiently introduced and fixed within a pull request, a scope more representative of the typical software development life cycle. Furthermore, we find that syntax errors are the most common failure type. This is surprisingly high for human-committed errors, given that modern IDEs typically provide real-time syntax checking. For LLM-generated errors, this result is less surprising, as current models are known to be prone to syntactical mistakes (tambon2025bugs).
Furthermore, we analyzed the statistical properties of the code changes, as shown in Table 3(b). We find that Augmented Dependency Errors involve substantially larger code changes. Our curation method for these errors selects commits with build file modifications, which often correlate with more substantial source code changes. As we demonstrate in our results (Table 8), the magnitude of the code change (i.e., files and lines modified) that led to the error is a reasonable proxy for the difficulty of the repair task, rather than the category of the error. A full distribution is in Appendix E.
@@ 3.1 Motivation: Providing the Right Tools
LLMs demonstrate strong proficiency with common shell commands and can often infer how to use novel tools from their descriptions (patil2025bfcl; zeng2025glm; tbench_2025). However, our preliminary analysis revealed a critical limitation when applying this general capability to a specialized domain like Android build repair. Although an LLM may possess the high-level knowledge to understand a build failure, it struggles to translate this knowledge into the correct sequence of low-level shell commands. For example, correctly synthesizing and sequencing a complex command like ./gradlew assembleDebug along with necessary environment variables is highly error-prone compared to the high-level decision of when to run a build.
We find this gap between high-level reasoning and low-level execution (Tables 4 & 6) likely occurs because while common shell commands (e.g., ls, grep) are prevalent in training data (liu2024toolace; patil2024gorilla), the multi-step sequences required for domains like Android builds are less common. An agent with a general shell often knows the relevant commands individually but fails to apply them in the correct context, leading to inefficient trial-and-error loops. This phenomenon is analogous to providing a specialist with a curated set of instruments rather than an exhaustive, unorganized toolkit.
Our approach addresses this through Tool Bridging, which replaces general-purpose tools with domain-aware abstractions. These specialized tools translate shell commands into high-level, API-like actions, presenting a format that LLMs handle more reliably and constraining the action space (patil2025bfcl). This design allows the model to focus its reasoning on debugging rather than command synthesis, bridging the gap between conceptual understanding and correct execution.
@@ 3.2 Execution Environment
All experiments were conducted on a standardized Linux machine (256 CPU cores, 512 GB RAM) installed with the Android SDK. To accommodate the diverse requirements of different projects, we installed the major versions of the Java Development Kit (JDK), including versions 11, 17, 20, 21, 22, and 23. Following the methodology of prior work (liu2024understanding), we build each application using the command ./gradlew assembleDebug --parallel. Building Android applications is computationally intensive, with build times for some projects extending up to 15 minutes on a 256-core machine. We containerized each build process within an isolated Jupyter Notebook (kluyver2016jupyter). This approach enabled us to run multiple tasks concurrently, significantly improving experimentation throughput. Additionally, to ensure a clean state between runs, we run ./gradlew clean --stop before attempting to build, preventing cached states from affecting the current build.
@@ 3.3 Approaches
We evaluate the effectiveness of an agent-based approach to repairing Android build errors by comparing several configurations designed to isolate the impact of domain-specific tooling. All of our agent configurations use Gemini-2.5-Pro (comanici2025gemini) as the core LLM.
As a baseline representing a state-of-the-art coding agent, we use the official, unmodified Gemini-CLI open-source implementation (google_gemini_cli). To allow for controlled experiments, we also developed a validated Python replica of the Gemini-CLI agent (Appendix B). This replica preserves the original agentic loop and prompting (Appendix C) but allows us to easily customize the available toolset. We use this replica for the Gemini-CLI (Read/Write Only) configuration and our proposed method, GradleFixer. We evaluate the following five distinct approaches:
  • Coding-Assistant: A baseline that simulates a user-in-the-loop assistant. The agent is provided with file context and an error log but can only propose code modifications. It cannot execute commands, browse the file system, or run the build. We use the open-source Aider (aider-website) framework for this configuration.
  • Hierarchical Agent: A two-agent framework where a primary LLM agent reasons about the problem and then delegates code editing tasks to the Coding-Assistant by invoking it as a tool, similar to ishibashi2024self; liu2024towards.
@@
  • Gemini-CLI (Read/Write and Shell): The standard open-source Gemini-CLI agent google_gemini_cli. We use the unmodified open-source version for this baseline. It is given file read/write capabilities and a general-purpose shell tool, allowing it to execute arbitrary commands.
@@ 3.4 GradleFixer Domain-Specific Tools
  • TOOL_A: [REDACTED].
  • TOOL_B: [REDACTED].
  • TOOL_C: [REDACTED].
All three tools are intentionally designed as wrappers for specialized shell commands. The outputs of these tools are not specially structured or processed. This design ensures the output is identical to that of a raw shell command, thereby isolating the impact of the tool invocation itself from any benefit of structured output. Full tool descriptions are available in Appendix D.
@@ 4 Results and Discussions
While the open-source Gemini-CLI already abstracts some shell commands (e.g., ls, grep) into dedicated tools, our method takes this direction further. We find that replacing the generic shell with the domain-specific tools TOOL_A, TOOL_B, and TOOL_C, leads to a dramatic performance improvement. The most substantial gain was on Dependency issues, which is where Gemini-CLI (Shell) struggled the most.
@@ 4.1 Impact of domain-specific tools
We study the impact of the domain-specific tools, which are all abstractions of shell commands. We evaluated agents equipped with different tool combinations, from the most general (shell) to the most specific (TOOL_A). The results are shown in Table 5. For ablations, LLMs are budgeted to 30 calls per task.
@@
  Only TOOL_B 55.8%
  Only TOOL_A 63.4%
@@
  TOOL_A + TOOL_B 69.7%
  TOOL_A + TOOL_C 69.8%
  Shell + TOOL_A + TOOL_B + TOOL_C 70.7%
  TOOL_A + TOOL_B + TOOL_C (Ours) 74.0%
Table 6: Tool usage frequency (%) across agent configurations.
@@
TOOL_C - 13.313.3* - - $0.6$
TOOL_B - 4.84.8* $24.7$ - $1.8$
TOOL_A - 20.720.7* - $20.2$ $14.8$
@@
Our results show a clear trend of improved performance as tools become more specific. The agent with only the shell tool improves over the baseline. However, performance increases with the more abstract TOOL_B tool and is highest with the most specific tool, TOOL_A. This result is particularly notable because shell is a superset of TOOL_B, which is a superset of TOOL_A. The agent is most effective when given the most constrained, highest-abstraction tool, suggesting the performance gain comes from focus, not just capability. Performance further improved as tools were combined, with our full method achieving the highest resolve rate and demonstrating the value of well-designed shell abstractions. Furthermore, our full method, which excludes the shell, outperforms the combination that includes it (74.0% vs 70.7%), suggesting that a constrained and relevant toolset is more effective than an unrestricted one. These observations align with findings in wang2023voyager, where specialized toolsets enable continuous performance improvement while more generic agents plateau.
@@
These results validate our Tool Bridging strategy, which bridges the gap between an LLM’s high-level knowledge and its low-level execution. The shell-based agent’s behavior suggests such a gap exists. Table 6 shows it frequently attempts the correct raw commands for building (20.7%) and changing Java versions (13.3%), yet its low pass rate (Table 4) demonstrates it struggles to apply them effectively in a raw shell. This pattern suggests the agent possesses the high-level knowledge of what operations are needed but struggles to correctly sequence and apply those operations. We illustrate these patterns in our case studies where the shell-based agent becomes trapped by misleading logs or reverts correct fixes (Appendix F).
@@
Tool Bridging addresses this gap by providing domain-aware abstractions. By replacing the general shell with specific tools like TOOL_A, we reframe complex shell interactions as simple, API-like calls. This aligns with the format LLMs are trained to use reliably (patil2025bfcl) and constrains the action space to relevant operations. This allows the model to leverage its high-level reasoning without failing on low-level command syntax and sequencing. Our results support this hypothesis. As shown in Table 5, performance improves as tools become more specific, suggesting LLMs know which commands are necessary but struggles to sequence them correctly as tools become less specific. GradleFixer achieves superior performance while calling TOOL_C far less frequently (0.6%), suggesting effective targeted tool use rather than exploratory use.
@@
In contrast to a general-purpose shell, a dedicated tool like TOOL_B could be priming the LLM to Android knowledge not only through its API-like structure but also through its name and description. This explicit contextual signal appears to help the model more reliably apply its latent knowledge of Android builds, steering it away from the incorrect interactions with a generic shell, as demonstrated in our case studies (Appendix F).
@@
REDACTED
We also include an evaluation using GPT-5-Mini to show that our finding is not limited to a single family of LLMs. While OpenAI provides its own coding agent, Codex openai-codex, it is not open-source, so its exact toolset cannot be reproduced. We instead use the same tools as Gemini-CLI. The results are shown in Table 7.
We find that providing a tool usage guidance prompt moderately improves performance over the baseline. This shows that guiding the LLM toward important tool use is an effective strategy. However, the prompted agent still performs significantly worse than our proposed agent, GradleFixer, which is equipped with domain-specific tools. This performance gap suggests that providing domain knowledge through dedicated tools is more effective than providing it through prompts for a general-purpose tool. We hypothesize this is because dedicated tools present operations in a more reliable, API-like format and constrain the action space, which more effectively connects the agent’s high-level reasoning to the low-level execution required for builds.

### v3

@@ Abstract
Ha Min Son1,2 Huan Ren2 Xin Liu1 Zhe Zhao1,* 1Department of Computer Science, University of California, Davis 2CodeDroid LLC
@@ 1 Introduction
Android is the largest and most active mobile development platform in the world, with a market share of 73.9% of the global mobile OS market, over 2.2 million apps on the Google Play Store, and over 1.4 million questions tagged "android" on Stack Overflow (Statista, 2024; StatCounter Global Stats, ; Stack Overflow, ). Despite its popularity, a practical challenge remains: applications cannot reliably be built. A comprehensive study of 5,222 apps revealed that only 31.3% could be built successfully out-of-the-box (Liu et al., 2024a). In our dataset, we find the primary source of issues is syntax errors (Table 3(a)). While Continuous Integration and Continuous Deployment (CI/CD) pipelines are a well-established practice for mitigating these errors (Rostami Mazrae et al., 2023; Baitha et al., 2024; Ghaleb et al., 2024; Liu et al., 2022), Liu et al. (2024a) reported only 21.8% out of the 5,222 apps had CI/CD scripts. Furthermore, even with CI/CD, maintaining a stable build requires significant effort. Hassan et al. (2017) found that over 74% of CI/CD configuration changes are dedicated to build fixing and environment updates. This maintenance is necessary to address a complex range of errors, including syntax mistakes, missing resource files, configuration errors, and library unavailability.
Large Language Models (LLMs) and their application in Automated Program Repair (APR) offer a promising alternative (Zhang et al., 2024a; Yin et al., 2024; Xia et al., 2023; Jin et al., 2023). However, the application of LLMs to Android build errors remains an underexplored area. Existing LLM coding benchmarks, like SWE-bench (Jimenez et al., 2023), primarily evaluate models on resolving bugs from GitHub issues, not the transient, environment-specific build failures common in Android development. Furthermore, existing datasets of Android build failures (Liu et al., 2024a; Hassan et al., 2017) capture a repository at a single, broken point in time, making fixes difficult to verify as feasible without a corresponding working build to serve as ground truth.
@@
Second, we propose GradleFixer, an agent-based approach for build error fixing using domain-specific tools. Inspired by prior work showing that specialized tools improve agent performance (Singh et al., 2025; Wang et al., 2023) (details in Sec. 5), GradleFixer equips an LLM agent with domain-specific tools, which are simple wrappers for shell commands, to inspect and manipulate the Gradle build environment. Our experiments show that GradleFixer significantly outperforms state-of-the-art baselines across all categories of build failures (Figure 1). We infer from GradleFixer’s success that LLMs have the high-level knowledge needed to fix build errors. The poorer performance of shell-based agents, however, suggests this knowledge fails to transfer to effective low-level execution. This gap explains why other effective agents also rely on specialized toolsets (Wang et al., 2023; Singh et al., 2025), although these studies do not explore this choice as a research question. Our approach, which we term Tool Bridging, addresses this gap by replacing a general-purpose shell with domain-aware abstractions to bridge high-level reasoning to correct low-level execution. We hypothesize this strategy improves performance through two simultaneous mechanisms. First, by providing API-like tools that abstract away complex command syntax, it allows the model to focus on reasoning rather than implementation. Second, it constrains the action space, reducing irrelevant operations. By reframing the task into a format the model uses more reliably (Liu et al., 2024b; Zhang et al., 2024b, 2025), our strategy connects the LLM’s high-level reasoning to effective low-level execution. Our contributions are summarized as follows:
@@ 2.1 Project Curation
AndroidBuildBench is curated from 43 active Android projects on GitHub. These were selected by filtering the top 100 most-starred applications for Java/Kotlin usage, >500 stars, >100 closed pull requests, recent commits, and a Gradle build script. This process yielded 43 open-source projects (Appendix F). A summary of the project statistics is shown in Table 1. We focus on the Gradle build system, as it is used by 95.1% of Android applications according to a recent study (Liu et al., 2024a).
@@ 2.3 Error Categorization
We manually analyzed the 184 test set issues, adopting the five build failure categories proposed by Liu et al. (2024a).
@@
Our analyses are presented in Table 3(a). A notable finding is the absence of NDK errors. We posit this absence results from our curation method, which is anchored to successfully built pull requests. NDK errors often result from build machine environment incompatibilities. Since our method requires a successful build as an endpoint, it naturally filters out states with systemic environment issues. The errors captured are those transiently introduced and fixed within a pull request, a scope more representative of the typical software development life cycle. Furthermore, we find that syntax errors are the most common failure type. This is surprisingly high for human-committed errors, given that modern IDEs typically provide real-time syntax checking. For LLM-generated errors, this result is less surprising, as current models are known to be prone to syntactical mistakes (Tambon et al., 2025).
Furthermore, we analyzed the statistical properties of the code changes, as shown in Table 3(b). We find that Augmented Dependency Errors involve substantially larger code changes. Our curation method for these errors selects commits with build file modifications, which often correlate with more substantial source code changes. As we demonstrate in our results (Table 8), the magnitude of the code change (i.e., files and lines modified) that led to the error is a reasonable proxy for the difficulty of the repair task, rather than the category of the error. A full distribution is in Appendix D.
@@ 3.1 Motivation: Providing the Right Tools
LLMs demonstrate strong proficiency with common shell commands and can often infer how to use novel tools from their descriptions (Patil et al., 2025; Zeng et al., 2025; Team, 2025). However, our preliminary analysis revealed a critical limitation when applying this general capability to a specialized domain like Android build repair. Although an LLM may possess the high-level knowledge to understand a build failure, it struggles to translate this knowledge into the correct sequence of low-level shell commands. For example, correctly synthesizing and sequencing a complex command like ./gradlew assembleDebug along with necessary environment variables is highly error-prone compared to the high-level decision of when to run a build.
We find this gap between high-level reasoning and low-level execution (Tables 4 & 6) likely occurs because while common shell commands (e.g., ls, grep) are prevalent in training data (Liu et al., 2024b; Patil et al., 2024), the multi-step sequences required for domains like Android builds are less common. An agent with a general shell often knows the relevant commands individually but fails to apply them in the correct context, leading to inefficient trial-and-error loops. This phenomenon is analogous to providing a specialist with a curated set of instruments rather than an exhaustive, unorganized toolkit.
Our approach addresses this through Tool Bridging, which replaces general-purpose tools with domain-aware abstractions. These specialized tools translate shell commands into high-level, API-like actions, presenting a format that LLMs handle more reliably and constraining the action space (Patil et al., 2025). This design allows the model to focus its reasoning on debugging rather than command synthesis, bridging the gap between conceptual understanding and correct execution.
@@ 3.2 Execution Environment
All experiments were conducted on a standardized Linux machine (256 CPU cores, 512 GB RAM) installed with the Android SDK. To accommodate the diverse requirements of different projects, we installed the major versions of the Java Development Kit (JDK), including versions 11, 17, 20, 21, 22, and 23. Following the methodology of prior work (Liu et al., 2024a), we build each application using the command ./gradlew assembleDebug --parallel. Building Android applications is computationally intensive, with build times for some projects extending up to 15 minutes on a 256-core machine. We containerized each build process within an isolated Jupyter Notebook (Kluyver et al., 2016). This approach enabled us to run multiple tasks concurrently, significantly improving experimentation throughput. Additionally, to ensure a clean state between runs, we run ./gradlew clean --stop before attempting to build, preventing cached states from affecting the current build.
@@ 3.3 Approaches
We evaluate the effectiveness of an agent-based approach to repairing Android build errors by comparing several configurations designed to isolate the impact of domain-specific tooling. All of our agent configurations use Gemini-2.5-Pro (Comanici et al., 2025) as the core LLM.
As a baseline representing a state-of-the-art coding agent, we use the official, unmodified Gemini-CLI open-source implementation (Google Cloud, ). To allow for controlled experiments, we also developed a validated Python replica of the Gemini-CLI agent (Appendix A). This replica preserves the original agentic loop and prompting (Appendix B) but allows us to easily customize the available toolset. We use this replica for the Gemini-CLI (Read/Write Only) configuration and our proposed method, GradleFixer. We evaluate the following five distinct approaches:
  • Coding-Assistant: A baseline that simulates a user-in-the-loop assistant. The agent is provided with file context and an error log but can only propose code modifications. It cannot execute commands, browse the file system, or run the build. We use the open-source Aider (Aider, ) framework for this configuration.
  • Hierarchical Agent: A two-agent framework where a primary LLM agent reasons about the problem and then delegates code editing tasks to the Coding-Assistant by invoking it as a tool, similar to Ishibashi and Nishimura (2024); Liu et al. (2024c).
@@
  • Gemini-CLI (Read/Write and Shell): The standard open-source Gemini-CLI agent Google Cloud . We use the unmodified open-source version for this baseline. It is given file read/write capabilities and a general-purpose shell tool, allowing it to execute arbitrary commands.
@@ 3.4 GradleFixer Domain-Specific Tools
  • run_build: Executes the standard build (./gradlew assembleDebug) and returns the raw output.
  • run_gradle: A more general tool that can execute any Gradle command (./gradlew <task>). This tool is a superset of run_build.
  • change_java_version: A dedicated function that switches the active JDK in the environment (e.g., export JAVA_HOME=...).
All three tools are intentionally designed as wrappers for specialized shell commands. The outputs of these tools are not specially structured or processed. This design ensures the output is identical to that of a raw shell command, thereby isolating the impact of the tool invocation itself from any benefit of structured output. Full tool descriptions are available in Appendix C.
@@ 4 Results and Discussions
While the open-source Gemini-CLI already abstracts some shell commands (e.g., ls, grep) into dedicated tools, our method takes this direction further. We find that replacing the generic shell with the domain-specific tools run_build, run_gradle, and change_java_version, leads to a dramatic performance improvement. The most substantial gain was on Dependency issues, which is where Gemini-CLI (Shell) struggled the most.
@@ 4.1 Impact of domain-specific tools
We study the impact of the domain-specific tools, which are all abstractions of shell commands. We evaluated agents equipped with different tool combinations, from the most general (shell) to the most specific (run_build). The results are shown in Table 5. For ablations, LLMs are budgeted to 30 calls per task.
@@
  Only Gradle (run_gradle) 55.8%
  Only build (run_build) 63.4%
@@
  Build + Gradle 69.7%
  Build + Java 69.8%
  Shell + Build + Gradle + Java 70.7%
  Build + Gradle + Java (Ours) 74.0%
Table 6: Tool usage frequency (%) across agent configurations. Agents are: Base (no shell), Shell (base with shell), Gradle (Only gradle), Build (Only Build), and Ours (Gradle, Build, and Java).
@@
Change Java - 13.313.3* - - $0.6$
Gradle - 4.84.8* $24.7$ - $1.8$
Build - 20.720.7* - $20.2$ $14.8$
@@
Our results show a clear trend of improved performance as tools become more specific. The agent with only the shell tool improves over the baseline. However, performance increases with the more abstract run_gradle tool and is highest with the most specific tool, run_build. This result is particularly notable because shell is a superset of run_gradle, which is a superset of run_build. The agent is most effective when given the most constrained, highest-abstraction tool, suggesting the performance gain comes from focus, not just capability. Performance further improved as tools were combined, with our full method (Build+Gradle+Java) achieving the highest resolve rate and demonstrating the value of well-designed shell abstractions. Furthermore, our full method, which excludes the shell, outperforms the combination that includes it (74.0% vs 70.7%), suggesting that a constrained and relevant toolset is more effective than an unrestricted one. These observations align with findings in Wang et al. (2023), where specialized toolsets enable continuous performance improvement while more generic agents plateau.
@@
These results validate our Tool Bridging strategy, which bridges the gap between an LLM’s high-level knowledge and its low-level execution. The shell-based agent’s behavior suggests such a gap exists. Table 6 shows it frequently attempts the correct raw commands for building (20.7%) and changing Java versions (13.3%), yet its low pass rate (Table 4) demonstrates it struggles to apply them effectively in a raw shell. This pattern suggests the agent possesses the high-level knowledge of what operations are needed but struggles to correctly sequence and apply those operations. We illustrate these patterns in our case studies where the shell-based agent becomes trapped by misleading logs or reverts correct fixes (Appendix E).
@@
Tool Bridging addresses this gap by providing domain-aware abstractions. By replacing the general shell with specific tools like run_build, we reframe complex shell interactions as simple, API-like calls. This aligns with the format LLMs are trained to use reliably (Patil et al., 2025) and constrains the action space to relevant operations. This allows the model to leverage its high-level reasoning without failing on low-level command syntax and sequencing. Our results support this hypothesis. As shown in Table 5, performance improves as tools become more specific, suggesting LLMs know which commands are necessary but struggles to sequence them correctly as tools become less specific. GradleFixer achieves superior performance while calling change_java_version far less frequently (0.6%), suggesting effective targeted tool use rather than exploratory use.
@@
In contrast to a general-purpose shell, a dedicated tool like run_gradle could be priming the LLM to Android knowledge not only through its API-like structure but also through its name and description. This explicit contextual signal appears to help the model more reliably apply its latent knowledge of Android builds, steering it away from the incorrect interactions with a generic shell, as demonstrated in our case studies (Appendix E).
@@
- You can change the Java version by setting the environment variable: export JAVA_HOME=.../[JAVA_VERSION]/lib/jvm
- You can run specific Gradle tasks using the shell command: ./gradlew [TASK_NAME]
- Once you apply a fix, you can test it using the shell command: ./gradlew assembleDebug
We also include an evaluation using GPT-5-Mini to show that our finding is not limited to a single family of LLMs. While OpenAI provides its own coding agent, Codex OpenAI (2025), it is not open-source, so its exact toolset cannot be reproduced. We instead use the same tools as Gemini-CLI. The results are shown in Table 7.
We find that providing a tool usage guidance prompt moderately improves performance over the baseline. This shows that guiding the LLM toward important tool use is an effective strategy. However, the prompted agent still performs significantly worse than our proposed agent, GradleFixer, which is equipped with domain-specific tools (Build, Gradle, and Java). This performance gap suggests that providing domain knowledge through dedicated tools is more effective than providing it through prompts for a general-purpose tool. We hypothesize this is because dedicated tools present operations in a more reliable, API-like format and constrain the action space, which more effectively connects the agent’s high-level reasoning to the low-level execution required for builds.